OpenStudy (anonymous):

If r=9, b=5, and g=-6, what does (r+b-g)(b+g) equal?

OpenStudy (anonymous):

You have to just plug in the values..

OpenStudy (anonymous):

\[\frac{r+b - g}{b + g} \implies \frac{9 +5 - (-6)}{5 + (-6)} = ??\]

OpenStudy (anonymous):

Remember: \(-(-) = +\) \(+(-) = -\)

OpenStudy (anonymous):

@waterineyes thank you so much! I got it from here!